OpenStudy (anonymous):

seldamat Solve the differential: y''+[(1/x)*(y')]-[(1/x^2)*y]=ln(x) What would be the best way to approach a problem like this? Variation of parameters?

OpenStudy (amistre64):

you stated from the last one:"I was thinking to solve, would be to write out x^2*yp''+x*yp'+yp = x^2*lnx as we solved for them - then equate the coefficients to the right hand side" much better now

OpenStudy (amistre64):

this would have been sufficient for constant coeffs i beleive

OpenStudy (anonymous):

yes, i am trying that now,

OpenStudy (anonymous):

nm .. lol sooo should I just plug in then what?

OpenStudy (amistre64):

but since our "coeffs" are functions, that wont be sufficient for our needs

OpenStudy (amistre64):

recall that we solced the homogenous part: and im going to forgo the latex since thats alot of the slowdown yh = c1 x^-1 + c2 x = 0

OpenStudy (amistre64):

since our coeffs are functions of x, and not constants; we let c1 and c2 equal arbitrary functions in x and equate this to the yp part

OpenStudy (anonymous):

yup! i follow

OpenStudy (amistre64):

yp = a(x) x^-1 + b(x) x yp' = -a(x) x^-2 + a'(x) x^-1 +b'(x) x + b(x)x' ----------------------------- -a(x) x^-2 + 0 + b(x) good so far?

OpenStudy (anonymous):

yes i have that

OpenStudy (amistre64):

yp'' = 2a(x) x^-3 - a'(x) x^-2 + b'(x) plug these values into the eq to simplify it such that it equates to x^2 ln(x)

OpenStudy (anonymous):

yup that is what im working on now

OpenStudy (anonymous):

yup so i get teh same answer= -A' + B'x^2 = x^2*ln(x)

OpenStudy (amistre64):

x^2 yp'' = 2a(x) x^-1 - a'(x) + b'(x) x^2 + x yp' = -a(x) x^-1 + b(x) x - yp = -a(x) x^-1 -b(x) x --------------------------------- x^2 ln(x) = - a'(x) + b'(x) x^2

OpenStudy (amistre64):

good, now we have 2 equations in A' and B' to solve for a system of eqs

OpenStudy (anonymous):

yh and the equation above?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

so yh = A'x^-1 + B'x = 0

OpenStudy (amistre64):

yes, and with any luck our longest version of this method is accurate lol for our yp in A' and B'

OpenStudy (anonymous):

sooo i got B' = ln(x) / 2 A' = (-ln(x)/2)^2

OpenStudy (anonymous):

er minus on the outside

OpenStudy (anonymous):

A' = -(ln(x)/2)^2 B' = ln(x)/2

OpenStudy (anonymous):

if that looks right im gonna try and solve for y

OpenStudy (amistre64):

you have to find A and B so integrate thos

OpenStudy (anonymous):

yup, just realized that as i started to plug in

OpenStudy (lalaly):

i can show u how to do it by variation of parameters if u want,, i think its easier in this case

OpenStudy (amistre64):

this is variation of paramenters; the long version

OpenStudy (amistre64):

he hasnt gotten to the wronskian method yet

OpenStudy (lalaly):

oh lol i saw the letters i assumed its undetermined coeffecients,, never mind:)

OpenStudy (anonymous):

the wronskian sounds easier (still working, solving for A now - had to brush up on integration by parts lol)

OpenStudy (amistre64):

i had to go to paper to keep track of things :) i get: a' = -x^2 ln(x) b' = ln(x)/(1-x)

OpenStudy (amistre64):

|a' = |-x^2 ln(x) = -ln(x) x^3/3 - | -x^3/3x = -ln(x) x^3/3 +| x^2/3 \(\large = \frac{x^3(1-3ln(x))}{9}\)

OpenStudy (amistre64):

the wolf says we end up with: y = c1 x^-1 + c2 x - ln(x)

OpenStudy (anonymous):

i got something different

OpenStudy (anonymous):

still simplifying though

OpenStudy (amistre64):

yeah, my mathing is off somplace i think

OpenStudy (amistre64):

the wronskian is way way better for this mess :/

OpenStudy (anonymous):

y = c1x^-1 + c2x + ln^2(x) - 2ln(x) + (x^2*ln(x)-x) / 2

OpenStudy (anonymous):

where A' = ln^2(x)/4 and B' = ln(x) / 2

OpenStudy (anonymous):

B = 1/2*(ln(x) - x) A = [ln^2(x)]*x - integrand(x*2ln(x)*1/x dx) (integration by parts) solve and you get A = ln^2(x)*x - 2*x(ln(x) -1 )

OpenStudy (amistre64):

dbl chk your system of eqs for a' and b'

OpenStudy (anonymous):

yea i'm pretty sure that they are right

OpenStudy (anonymous):

A' and B' that is

OpenStudy (amistre64):

\[\begin{align}x^{-1}A'+xB' &= 0\\-x^{-1}A'+1B' &= ln(x)\end{align}\]

OpenStudy (amistre64):

(1+x)B' = ln(x) B' = ln(x)/(1+x)

OpenStudy (amistre64):

x^-1A' = -xB' A' = -x^2 B'

OpenStudy (anonymous):

how did you get -A'/x + B = ln(x)

OpenStudy (amistre64):

that was the end results of the plugging and playing with the yp and its derivatives

OpenStudy (anonymous):

[-A'+2A/x + B'x^2] + [-A/x + B/x] - [A/x + Bx] = x^2*ln(x)

OpenStudy (amistre64):

-A' x + B'x^2 = x^2 ln(x) ; divide off the x^2 -A'/x + B' = ln(x)

OpenStudy (anonymous):

then -A' + 2A/x - 2A/x + B'x^2 + Bx - Bx = x^2 * ln(x)

OpenStudy (amistre64):

-A'x at the start, see your typo

OpenStudy (anonymous):

the first expression i sent is after you distribute the x^2 - once you do that, you lose the x on A'

OpenStudy (amistre64):

x/x^2 = 1/x

OpenStudy (amistre64):

-A' x/x^2 = -A/x

OpenStudy (anonymous):

x^2(yp'') = x^2 ( 2Ax^-3 - x^-2A' + B')

OpenStudy (amistre64):

yeah, i just seen that lol

OpenStudy (amistre64):

prolly why my results were so ugly ...

OpenStudy (anonymous):

haha its ok, my result is ugly too

OpenStudy (amistre64):

a'/x + b' x = 0 <-- * -/x -a'/x^2 + b' = ln(x) -a'/x^2 - b' x = 0 -a'/x^2 + b' = ln(x) -------------------- -2a'/x^2 = ln(x) a' = -ln(x)/2 -ln(x)/2x^2 - b' x = 0 -ln(x)/2x - b' = 0 -ln(x)/2x^3 = b' now whatd i miss on that :)

OpenStudy (amistre64):

other than the x^2 .....

OpenStudy (anonymous):

a' = -ln(x)/2 * x^2 right?

OpenStudy (amistre64):

a' = -x^2 ln(x)/2 -x^2 ln(x)/2x + b' x = 0 : /x -x^2 ln(x)/2x^2 + b' = 0 -ln(x)/2 + b' = 0 b' = ln(x)/2 .....maybe?

OpenStudy (amistre64):

i have a theory; since i had to make room for the wronskian method; i had to forget alot of algebra in the process :)

OpenStudy (anonymous):

i suck at alegbra... lol so I should have tons of room for the wronskian i hope (according to your theory :P )

OpenStudy (anonymous):

i need to eat some lunch.. im out of brain juice, then ill com eback to check work and see if there are any obvious errors

OpenStudy (anonymous):

i looked at the wolfram solution - it has imaginary numbers in it ...

OpenStudy (anonymous):

ok brb before i starve myself

OpenStudy (amistre64):

|b'= | ln(x)/2 = x ln(x)/2 - | x/2x = x ln(x)/2 - x/2 |a' =|-x^2 ln(x)/2 = -x^3 ln(x)/6 + | x^3/6x = -x^3 ln(x)/6 + x^3/18 a/x = -x^2 ln(x)/6 + x^2/18 +bx = x^2 ln(x)/2 - x^2/2 ----------------------------- yp = x^2/2 (1-3ln(x)+9ln(x)-9) yp = x^2/2 (6ln(x)-8) = x^2 (3ln(x)-4) but typing and mathing fractions and such aint my forte so youd better double chk it :)

OpenStudy (anonymous):

I checked it btw - and it worked like a charm. Thank you so much! I've learned a lot you know.